OpenStudy (anonymous):

please help What are the vertices, foci and asymptotes of the hyperbola with equation 16x^2-4y^2=64

OpenStudy (anonymous):

V (2,0) & (-2,0) Co-V (0,4) & (0,-4) foci (3,0) & (-3,0) asymptotes y=-1/2x and y = 1/2x

OpenStudy (anonymous):

how did you find them?

OpenStudy (anonymous):

oh your equation is x^2/4 - y^2/16 = 1 when you divide both sides with 64.then you find the a and a is the denominator before the subtraction sign. so a is 2, b is the denom. after the subtraction sign so its 4. then you use a^2 + b^2 = c^2 to find c which is +/_ 3

OpenStudy (anonymous):

then since a is under the x it means its a horizontal hyperbola. center is (0,0) so for the V add +/- 2 to the x leaving you with V(+/-2,0) then add the 4 to the y in (0,0) -> (0, +/-3) as your co vertices. also since your focus has to be on the same line as your V you add your +/-3 to the X of (0,0) leaving you with (+/-3,0) . then you can find your asymptote by using the formula y=+/- a/bx (for horizontal only)

OpenStudy (anonymous):

was that helpful

OpenStudy (anonymous):

@NataliaK Yes thank you!! :)

OpenStudy (anonymous):

:) glad I could help